# Metropolis-Hastings Sampling for 3D Gaussian Reconstruction

### v1

@@ Abstract
We propose an adaptive sampling framework for 3D Gaussian Splatting (3DGS) that leverages comprehensive multi-view photometric error signals within a unified Metropolis-Hastings approach. Traditional 3DGS methods heavily rely on heuristic-based density-control mechanisms (e.g., cloning, splitting, and pruning), which can lead to redundant computations or the premature removal of beneficial Gaussians. Our framework overcomes these limitations by reformulating densification and pruning as a probabilistic sampling process, dynamically inserting and relocating Gaussians based on aggregated multi-view errors and opacity scores. Guided by Bayesian acceptance tests derived from these error-based importance scores, our method substantially reduces reliance on heuristics, offers greater flexibility, and adaptively infers Gaussian distributions without requiring predefined scene complexity. Experiments on benchmark datasets, including Mip-NeRF360, Tanks and Temples, and Deep Blending, show that our approach reduces the number of Gaussians needed, enhancing computational efficiency while matching or modestly surpassing the view-synthesis quality of state-of-the-art models. Our project page is available at https://hjhyunjinkim.github.io/MH-3DGS.
@@ 1 Introduction
Novel view synthesis (NVS) focuses on creating photo-realistic images of a 3D scene from unseen viewpoints and is significant due to its broad applicability across various real-world scenarios. While Neural Radiance Fields (NeRF) [34] has largely advanced Novel View Synthesis by mapping the 3D location and view directions to a view-dependent color and volumetric density, its slow rendering hinders real-world deployment. To address these limitations, 3D Gaussian Splatting (3DGS) [18] recently emerged, enabling real-time photo-realistic rendering by representing complex scenes with explicit 3D Gaussians, each of which is projected to the screen and defined by its position, anisotropic covariance, and opacity.
@@
Despite these advantages, 3DGS and its extensions [18, 55, 16] rely heavily on heuristic-based adaptive density control mechanisms, such as cloning, splitting, and pruning by introducing fixed thresholds on opacity, size, and positional gradients in the view space. This leads to inflation in memory with redundant Gaussians when thresholds are loose or sacrifice fidelity when they are tight. To tackle this problem, Kheradmand et al. [20] recast 3DGS updates as Stochastic Gradient Langevin Dynamics (SGLD), framing Gaussian updates as Markov Chain Monte Carlo (MCMC) samples from an underlying probability distribution representing scene fidelity. As SGLD updates Gaussians in the direction of the image-reconstruction gradient and adds a small noise term, the procedure behaves like a gradient-based optimizer. Although it replaces heuristics with state transitions, it still fixes the total number of Gaussians upfront, limiting adaptability to scenes of varying complexity.
@@
Motivated by these considerations, we propose an adaptive sampling framework that uniquely leverages comprehensive multi-view error signals within a unified Metropolis-Hastings (MH) [33, 7] approach. We reformulate 3D Gaussian Splatting (3DGS) as an MH sampling process and newly derive a concise, rigorously proven acceptance probability grounded in the classical MH rule. By casting the densification and pruning stages of 3DGS as a probabilistic sampling problem, our method selectively adds Gaussians in regions with high visual or geometric significance as indicated by aggregated view-consistent photometric discrepancies. By deriving Gaussian-level importance scores from opacity and multi-view error metrics, we generate Gaussian candidates through coarse-to-fine proposals, which are subsequently accepted or rejected via Bayesian acceptance tests. This probabilistic process adaptively converges to a well-balanced Gaussian distribution with only minimal reliance on heuristics, requiring no prior knowledge of scene complexity while seamlessly integrating with existing 3DGS workflows to achieve superior reconstruction quality, capturing both coarse structures and fine details.
@@
We demonstrate the effectiveness of our approach on real-world scenes from the Mip-NeRF360 [1], Tanks and Temples [21], and Deep Blending [15] datasets. Our method reduces the number of Gaussians needed to represent each scene solely through an improved sampling mechanism, while consistently matching or surpassing the view-synthesis quality of 3DGS across diverse scenes. It also performs on par with, or outperforms Kheradmand et al. [20]. By coupling principled probabilistic inference with dense multi-view error signals, our method matches the rendering accuracy of 3DGS-MCMC [20] with a reduced number of Gaussians, demonstrating that comparable quality can be achieved with a leaner, more generalizable representation.
@@
  • We show through our experiments that our framework attains state-of-the-art view-synthesis quality with faster convergence.
@@ 2 Related Work
2.1 Neural Radiance Fields
Novel View Synthesis generates realistic images from novel viewpoints that differ from the original captures [55]. Neural Radiance Fields (NeRF) [34] have been widely explored in this field because of their remarkable ability to render photorealistic images: an MLP models the 3D scene from multi-view 2D images and enables novel-view generation with multi-view consistency, inspiring extensions that scale to large scenes[13, 50, 44], handle dynamic content [39, 40, 8, 47, 14], generalize to unseen scenes [37, 54], or dispense with known camera poses [31, 17, 25, 32, 2, 56]. Yet the significant limitation of NeRF is that novel-view synthesis demands extensive training and rendering time, prompting many works to accelerate these stages [43, 10, 35, 5, 41, 11].
Building on these advancements, recent research on NeRF explored efficient sampling and optimization to handle complex scenarios. Goli et al. [12] applies the Metropolis-Hastings algorithm to iteratively refine samples iteratively, accurately aligning partially overlapping scenes. Similarly, Bortolon et al. [3] integrates Metropolis-Hastings sampling to select key surface points for efficient ray casting, enabling real-time 6DoF pose estimation without an initial pose. While these methods primarily use Metropolis–Hastings for surface point selection, our work reformulates the entire 3D reconstruction problem as a Metropolis–Hastings sampling process. This perspective allows us to draw Gaussian samples without heuristic assumptions and further underscores the benefits of probabilistic sampling.
2.2 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [18] recently emerged as a compelling alternative to NeRF by using differentiable rasterization rather than relying on neural networks for volume rendering. 3DGS uses 3D Gaussian primitives to represent scenes, which are initialized from sparse point clouds generated by Structure-from-Motion (SfM) [42]. This representation enables efficient optimization through interleaved updates of each Gaussian’s position, covariances, opacities, and colors, represented by Spherical Harmonics.
Given 3DGS’s remarkable efficiency and high-resolution output, numerous studies have extended its application across a range of domains, from dynamic scene reconstruction [51, 29, 24, 49, 23], 3D generation [45, 53, 46, 57, 6, 27] to large-scale scene reconstruction [19, 26, 28]. A critical drawback, however, is that each Gaussian adds a significant memory burden due to attributes such as covariance matrices and spherical harmonics (SH) parameters, prompting several memory-efficient variants [22, 9, 36, 48, 38]. Yet these extensions depend on heuristic density-control rules, which constrain efficiency and generality. Addressing this limitation, we introduce a probabilistic framework that replaces fixed thresholds with a unified Metropolis-Hastings sampling strategy, absorbing both densification and pruning in 3DGS and thereby enhancing model efficiency and versatility.
@@ 2.3 Adaptive Density Control for 3D Gaussian Splatting
Kheradmand et al. [20] recently introduced the use of Stochastic Gradient Langevin Dynamics (SGLD), a modification of stochastic gradient descent and MCMC, for sampling and optimization in 3DGS, combined with an opacity-driven relocation strategy to shift low-opacity Gaussians towards regions with higher-opacity Gaussians, while enforcing a fixed limit on the total number of points. Additionally, Bulò et al. [4] introduces an error-based densification method that prioritizes areas with high perceptual errors, corrects opacity bias in cloning, and sets a global limit on the number of Gaussians. Similarly, Ye et al. [52] addresses the limitations of gradient collision in adaptive density control by proposing a homodirectional gradient as a criterion for densification. Moreover, Mallick et al. [30] presents a score-based guided densification approach that uses training-time priors to restrict uncontrolled model-size growth. As 3DGS-MCMC [20] is most closely aligned with our objectives and delivers the best overall reconstruction quality, it is our primary baseline for comparison.
In contrast, our proposed framework leverages a unified Metropolis-Hastings sampling strategy grounded in comprehensive multi-view error signals. Rather than relying on thresholds, our method computes Gaussian importance scores from aggregated multi-view photometric errors, dynamically adapting both the number and spatial distribution of Gaussians based on probabilistic inference. We derive the MH acceptance rule in the context of 3DGS and prove that our importance-weighted proposals preserve detailed balance with respect to the multi-view photometric likelihood. This rigorous formulation clearly distinguishes our framework from prior methods addressing this task. Consequently, our approach naturally scales with scene complexity, improving flexibility and efficiency.
@@ 3 Preliminary
where the covariance matrix is factorized into a scaling matrix $S$ and a rotation matrix $R$. To optimize the parameters of the 3D Gaussians for an accurate scene representation, it is essential to render them into images using a differentiable process. The transformed covariance matrix $\Sigma^{\prime}$, now in 2D, is calculated by the viewing transform $W$ and the Jacobian of the affine transformation of the projective transformation $J$ as $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$. The color $C$ of a pixel can be computed by blending $\mathcal{N}$ ordered points overlapping the pixel as $C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$, where $c_{i}$ and $\alpha_{i}$ represent the view-dependent color and the opacity of the point. These values are computed from a 2D Gaussian distribution with a covariance matrix $\Sigma$, scaled by a per-point optimized 3D Gaussian opacity.
Furthermore, to achieve a high-quality representation, 3DGS employs adaptive density control, a heuristic-rule-based strategy that fills under-reconstructed areas by cloning Gaussians and enhances over-reconstructed regions by dividing larger Gaussians. During each refinement step, 3DGS evaluates Gaussians based on their position, opacity, and size, removing those with low opacity or excessive size. For densification, 3DGS assesses the positional gradient $\nabla_{p}\mathcal{L}$ of each Gaussian, refining further if the gradient exceeds a predefined threshold $\tau_{p}$. If a Gaussian surpasses the size threshold $\tau_{s}$, indicating over-reconstruction, it is split into smaller Gaussians. Alternatively, under-reconstructed areas prompt Gaussian cloning by shifting along the gradient direction, effectively expanding Gaussian coverage for better scene representation.
@@ 4 Method
Our method begins with a key design choice: unlike 3DGS, which densifies points once a fixed threshold is reached, we employ a lean point-based sampler that keeps the most informative Gaussians, sharply reducing storage while preserving high-fidelity reconstructions. Sec.4.1 lays the theoretical groundwork for this: we extend the classical Metropolis–Hastings algorithm to 3DGS and rigorously derive scene-adaptive equations that generalize the traditional formulation. Sec.4.2 describes how the sampler proposes candidates guided by multi-view photometric and opacity errors, and Sec.4.3 presents the MH acceptance test, where we newly derive and prove a concise, practical expression for the MH acceptance probability. The complete algorithmic flow is summarized in Alg.1 and 2 in Appendix D.
@@ 4.1 Formulation of the Metropolis-Hastings Framework
Our main objective is eliminating heuristic dependence in adaptive density control by recasting densification and pruning as a principled sampling problem. Specifically, we interpret each modification of the Gaussian set as a proposal in a Metropolis-Hastings Markov Chain. (See Appendix F for the preliminary on the Metropolis-Hastings algorithm) Each step in the chain perturbs the current scene representation $\Theta$ by (i)inserting new Gaussians near pixels that still exhibit high multi–view error or (ii)relocating Gaussians whose opacity has collapsed. The proposed scene $\Theta^{\prime}$ is then accepted or rejected according to a probabilistic rule that, in expectation, favors configurations that reproduce the captured images better while remaining spatially sparse.
@@
where $c_{\Theta}(v)$ is the number of Gaussians in voxel $v$. This prior encourages at most one or two Gaussians per voxel. Since the cost grows logarithmically, empty cells get almost no penalty, but the penalty rises quickly once the voxel becomes crowded. Combining Eq. 2 and 4, we are able to define a negative log-posterior:
@@
where $Z$ is a normalization constant (See Appendix A for details). Our framework samples Gaussian configurations $\Theta$ in proportion to $\pi(\Theta)$, adaptively refining the 3D Gaussian representation to cover undersampled regions. To facilitate this sampling, we construct a Markov Chain whose states are successive scene representations $\Theta_{0},\Theta_{1},\dots,\Theta_{t}$. From a given state $\Theta_{t}$, we draw a proposal $\Theta^{\prime}$ by adding Gaussians in high-error areas or relocating low opacity Gaussians, with probability density $q(\Theta^{\prime}\mid\Theta_{t})$. The construction of $q(\Theta^{\prime}\!\mid\!\Theta_{t})$ is detailed in Section 4.2.
@@
where $\rho(\Theta^{\prime}\mid\Theta_{t})$ indicates the standard MH acceptance probability. At each step we draw $u\sim\mathcal{U}(0,1)$; if $u<\rho$, the proposal is accepted and we set $\Theta_{t+1}=\Theta^{\prime}$, otherwise we keep $\Theta_{t+1}=\Theta_{t}$. We explain how the MH rule accepts or rejects $\Theta^{\prime}$ again in Section 4.3. Over many iterations, this Markov Chain converges toward our target distribution $\pi(\Theta)$, systematically densifying the Gaussian set where it reduces loss and relocating Gaussians that fail to improve image fidelity.

### v2

@@ Abstract
We propose an adaptive sampling framework for 3D Gaussian Splatting (3DGS) that leverages comprehensive multi-view photometric error signals within a unified Metropolis-Hastings approach. Vanilla 3DGS heavily relies on heuristic-based density-control mechanisms (e.g., cloning, splitting, and pruning), which can lead to redundant computations or premature removal of beneficial Gaussians. Our framework overcomes these limitations by reformulating densification and pruning as a probabilistic sampling process, dynamically inserting and relocating Gaussians based on aggregated multi-view errors and opacity scores. Guided by Bayesian acceptance tests derived from these error-based importance scores, our method substantially reduces reliance on heuristics, offers greater flexibility, and adaptively infers Gaussian distributions without requiring predefined scene complexity. Experiments on benchmark datasets, including Mip-NeRF360, Tanks and Temples and Deep Blending, show that our approach reduces the number of Gaussians needed, achieving faster convergence while matching or modestly surpassing the view-synthesis quality of state-of-the-art models. Our project page is available at https://hjhyunjinkim.github.io/MH-3DGS.
@@ 1 Introduction
Novel view synthesis (NVS) focuses on creating photo-realistic images of a 3D scene from unseen viewpoints. It is significant due to its broad applicability across various real-world scenarios. While Neural Radiance Fields (NeRF) [36] has largely advanced Novel View Synthesis by mapping 3D locations and view directions to view-dependent colors and volumetric densities, its slow rendering speed hinders real-world deployment. To address these limitations, 3D Gaussian Splatting (3DGS) [20] recently emerged, enabling real-time photo-realistic rendering by representing complex scenes with explicit 3D Gaussians, each of which is projected to the screen and defined by its position, anisotropic covariance, and opacity.
@@
Despite these advantages, 3DGS and its extensions [18, 20, 56] rely heavily on heuristic-based adaptive density control mechanisms, such as cloning, splitting, and pruning by introducing fixed thresholds on opacity, size, and positional gradients in the view space. This leads to inflation of memory with redundant Gaussians when thresholds are loose, or to a sacrifice of fidelity when they are tight. To tackle this problem, Kheradmand et al. [22] recast 3DGS updates as Stochastic Gradient Langevin Dynamics (SGLD), framing Gaussian updates as Markov Chain Monte Carlo (MCMC) samples from an underlying probability distribution representing scene fidelity. As SGLD updates Gaussians in the direction of the image-reconstruction gradient and adds a small noise term, the procedure behaves like a gradient-based optimizer. Although it replaces heuristics with state transitions, it still fixes the total number of Gaussians upfront, limiting adaptability to scenes of varying complexity.
@@
Motivated by these considerations, we propose an adaptive sampling framework that uniquely leverages comprehensive multi-view error signals within a unified Metropolis-Hastings (MH) [6, 35] approach. We reformulate 3D Gaussian Splatting (3DGS) as an MH sampling process and newly derive a concise, rigorously proven acceptance probability grounded in the classical MH rule. By casting the densification and pruning stages of 3DGS as a probabilistic sampling problem, our method selectively adds Gaussians in regions with high visual or geometric significance as indicated by aggregated view-consistent photometric discrepancies. By deriving Gaussian-level importance scores from opacity and multi-view error metrics, we generate Gaussian candidates through coarse-to-fine proposals, which are subsequently accepted or rejected via Bayesian acceptance tests. This probabilistic process adaptively converges to a well-balanced Gaussian distribution with only minimal reliance on heuristics. It requires no prior knowledge of scene complexity while seamlessly integrating with existing 3DGS workflows to achieve better reconstruction quality, capturing both coarse structures and fine details.
@@
We demonstrate the effectiveness of our approach on real-world scenes from the Mip-NeRF360 [1], Tanks and Temples [23], and Deep Blending [17] datasets. Our method reduces the number of Gaussians needed to represent each scene solely through an improved sampling mechanism, while consistently matching or surpassing the view-synthesis quality of 3DGS across diverse scenes. Additionally, our approach converges faster than 3DGS-MCMC [22], achieving target PSNR values in fewer training iterations and reduced wall-clock time. By coupling principled probabilistic inference with dense multi-view error signals, our method matches the rendering accuracy of 3DGS-MCMC [22] with fewer Gaussians, demonstrating that comparable quality can be achieved with a leaner, more generalizable representation that converges more rapidly.
@@
  • Our experiments show that our framework achieves compelling view synthesis quality while converging faster.
@@ 2 Related Work
2.1 Neural 3D Scene Representation
Novel View Synthesis generates realistic images from novel viewpoints that differ from the original captures [56]. Neural Radiance Fields (NeRF) [36] pioneered this field by using MLPs to model 3D scenes from multi-view 2D images, inspiring extensions for large scenes [13, 46, 52], dynamic content [8, 14, 40, 41, 49], and pose-free reconstruction [2, 19, 27, 33, 34, 57]. However, NeRF’s extensive training and rendering time prompted numerous acceleration efforts [4, 10, 11, 37, 42, 45].
3D Gaussian Splatting (3DGS) [20] emerged as a compelling alternative, using differentiable rasterization of 3D Gaussian primitives initialized from sparse point clouds generated from Structure-from-Motion [44]. This approach enables efficient optimization and high-resolution output, spurring extensions in dynamic scenes [25, 26, 31, 51, 53], 3D generation [5, 29, 47, 48, 55], and large-scale reconstruction [21, 28, 30]. Yet, 3DGS suffers from a significant memory burden due to covariance matrices and spherical harmonics, leading to memory-efficient variants that rely on heuristic density-control rules [9, 15, 16, 24, 38, 39, 50].
Recent work has explored probabilistic sampling in 3D reconstruction: Goli et al. [12] and Bortolon et al. [3] apply Metropolis-Hastings for sample refinement and surface point selection in NeRF. While these methods use Metropolis-Hastings for specific components, our work reformulates the entire 3D reconstruction problem as a Metropolis-Hastings sampling process, replacing heuristic density controls in 3DGS with a unified probabilistic framework that enhances both efficiency and generality.
@@ 2.2 Adaptive Density Control for 3D Gaussian Splatting
Kheradmand et al. [22] recently introduced the use of Stochastic Gradient Langevin Dynamics (SGLD) for sampling and optimization in 3DGS, combined with an opacity-driven relocation strategy to shift low-opacity Gaussians towards higher-opacity regions, while enforcing a fixed limit on the total number of Gaussians. Additionally, Rota Bulò et al. [43] introduces error-based densification with perceptual error prioritization and sets a global limit on the number of Gaussians. Similarly, Ye et al. [54] addresses the limitations of gradient collision in adaptive density control by proposing a homodirectional gradient as a criterion for densification. Moreover, Mallick et al. [32] presents a score-based guided densification approach that uses training-time priors to restrict growth. More recently, Deng et al. [7] introduced deterministic geometric splitting rules for efficient density control. As 3DGS-MCMC [22] aligns most closely with our objectives and delivers the best overall reconstruction quality, it serves as our primary baseline.
In contrast, our proposed framework leverages a unified Metropolis-Hastings sampling strategy grounded in comprehensive multi-view error signals. Rather than relying on thresholds, our method computes Gaussian importance scores from aggregated multi-view photometric errors and dynamically adapts the number and distribution of Gaussians via probabilistic inference. We derive the MH acceptance rule in the context of 3DGS and prove that our importance-weighted proposals preserve detailed balance with respect to the multi-view photometric likelihood. This rigorous formulation distinguishes our framework from prior methods and enables natural scaling with scene complexity.
@@ 3 Preliminary
where the covariance matrix is factorized into a scaling matrix $S$ and a rotation matrix $R$. The transformed 2D covariance matrix $\Sigma^{\prime}$ is calculated by the viewing transform $W$ and the Jacobian of the affine transformation of the projective transformation $J$ as $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$. The color $C$ of a pixel can be computed by blending $\mathcal{N}$ ordered points overlapping the pixel as $C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$, where $c_{i}$ and $\alpha_{i}$ represent the view-dependent color and opacity computed from a 2D Gaussian distribution.
To achieve a high-quality representation, 3DGS employs adaptive density control that clones Gaussians in under-reconstructed areas and splits Gaussians in over-reconstructed regions. During densification, 3DGS evaluates the positional gradient $\nabla_{p}\mathcal{L}$ of each Gaussian, refining further if the gradient exceeds a predefined threshold $\tau_{p}$. If a Gaussian surpasses the size threshold $\tau_{s}$, indicating over-reconstruction, it is split into smaller Gaussians. Alternatively, under-reconstructed areas prompt cloning by shifting Gaussians along the gradient direction for better scene representation.
3.2 Metropolis-Hastings Algorithm
The Metropolis-Hastings algorithm [6, 35] is a widely used Markov Chain Monte Carlo (MCMC) method for sampling from complex, high-dimensional probability distributions. The objective is to construct a Markov chain whose stationary distribution $\pi(\mathbf{x})$ matches the target distribution, enabling efficient sampling where direct methods are impractical.
Given a target distribution $\pi(\mathbf{x})$ defined over a state space $\mathcal{X}$, the algorithm generates a sequence $\{\mathbf{x}^{(t)}\}_{t=1}^{N}$ of samples by iteratively proposing candidates from a proposal distribution $q(\mathbf{x},\mathbf{y})$ and accepting or rejecting them based on an acceptance probability. At each iteration $t$, a new candidate $\mathbf{y}$ is proposed from the distribution $q(\mathbf{x}^{(t)},\cdot)$. The acceptance probability is further computed as
@@ 4 Method
Our method begins with a key design choice: unlike 3DGS, which densifies points at fixed thresholds, we employ a lean point-based sampler that keeps the most informative Gaussians, sharply reducing storage while preserving fidelity. Section 4.1 lays the theoretical groundwork for this: we extend the Metropolis–Hastings algorithm to 3DGS and rigorously derive scene-adaptive equations that generalize the traditional formulation. Section 4.2 describes how the sampler proposes candidates guided by multi-view photometric and opacity errors, and Sec. 4.3 presents the MH acceptance test with a newly derived practical expression for the MH acceptance probability. The complete algorithmic flow is in Algorithm 1 and 2 in Appendix D.
@@ 4.1 Formulation of the Metropolis-Hastings Framework
Our main objective is to eliminate heuristic dependence in adaptive density control by recasting densification and pruning as a principled sampling problem. Specifically, we interpret each modification of the Gaussian set as a proposal in a Metropolis-Hastings Markov Chain. Each step in the chain perturbs the current scene representation $\Theta$ by (i)inserting new Gaussians near pixels that still exhibit high multi–view error or (ii)relocating Gaussians whose opacity has collapsed. The proposed scene $\Theta^{\prime}$ is then accepted or rejected according to a probabilistic rule that, in expectation, favors configurations that reproduce the captured images better while remaining spatially sparse.
@@
where $c_{\Theta}(v)$ is the number of Gaussians in voxel $v$. This prior encourages at most one or two Gaussians per voxel. Since the cost grows logarithmically, empty cells get almost no penalty, but the penalty rises quickly once the voxel becomes crowded. Combining Eq. 4 and 6, we are able to define a negative log-posterior:
@@
where $Z$ is a normalization constant (See Appendix A for details). Our framework samples Gaussian configurations $\Theta$ in proportion to $\pi(\Theta)$ and adaptively refines the 3D Gaussian representation to cover undersampled regions. To facilitate this sampling, we construct a Markov Chain whose states are successive scene representations $\Theta_{0},\Theta_{1},\dots,\Theta_{t}$. From a given state $\Theta_{t}$, we draw a proposal $\Theta^{\prime}$ by adding Gaussians in high-error areas or relocating low opacity Gaussians, with probability density $q(\Theta^{\prime}\mid\Theta_{t})$. The construction of $q(\Theta^{\prime}\!\mid\!\Theta_{t})$ is detailed in Sec. 4.2.
@@
where $\rho(\Theta^{\prime}\mid\Theta_{t})$ indicates the standard MH acceptance probability. At each step we draw $u\sim\mathcal{U}(0,1)$; if $u<\rho$, the proposal is accepted and we set $\Theta_{t+1}=\Theta^{\prime}$, otherwise we keep $\Theta_{t+1}=\Theta_{t}$. We explain how the MH rule accepts or rejects $\Theta^{\prime}$ again in Sec. 4.3. Over many iterations, this Markov Chain converges toward our target distribution $\pi(\Theta)$, systematically densifying the Gaussian set, reducing loss, and relocating Gaussians that fail to improve image fidelity.